### v1

# Intent Laundering: AI Safety Datasets Are Not What They Seem

Abstract
We systematically evaluate the quality of widely used AI safety datasets from two perspectives: in isolation and in practice. In isolation, we examine how well these datasets reflect real-world attacks based on three key properties: driven by ulterior intent, well-crafted, and out-of-distribution. We find that these datasets overrely on “triggering cues”: words or phrases with overt negative/sensitive connotations that are intended to trigger safety mechanisms explicitly, which is unrealistic compared to real-world attacks. In practice, we evaluate whether these datasets genuinely measure safety risks or merely provoke refusals through triggering cues. To explore this, we introduce “intent laundering”: a procedure that abstracts away triggering cues from attacks (data points) while strictly preserving their malicious intent and all relevant details. Our results indicate that current AI safety datasets fail to faithfully represent real-world attacks due to their overreliance on triggering cues. In fact, once these cues are removed, all previously evaluated “reasonably safe” models become unsafe, including Gemini 3 Pro and Claude Sonnet 3.7. Moreover, when intent laundering is adapted as a jailbreaking technique, it consistently achieves high attack success rates, ranging from 90% to over 98%, under fully black-box access. Overall, our findings expose a significant disconnect between how model safety is evaluated and how real-world adversaries behave.
Warning: This paper may contain AI-generated sensitive content.
1 Introduction
Safety alignment and safety datasets are the two pillars of AI safety (Beyer et al., 2025). Safety alignment focuses on post-training techniques that fortify models against adversarial attacks (Askell et al., 2021; Ouyang et al., 2022), while safety datasets serve to evaluate the robustness of these defenses (Mazeika et al., 2024; Zou et al., 2023). The credibility of these evaluations, therefore, depends largely on the quality of the safety datasets. As with other evaluation tasks, dataset quality is determined by how well these datasets represent real-world scenarios. However, unlike other datasets that approximate such scenarios through common patterns (e.g., math problems), in the context of AI safety, the real-world scenarios involve ulterior, well-crafted, and out-of-distribution attacks that datasets must capture. This makes the design and development of safety datasets fundamentally different from other datasets.
Figure 1: Overview of our intent laundering framework. Without the feedback loop, the framework illustrates the intent laundering procedure. With the loop, it shows how intent laundering operates as a jailbreaking technique. The process begins by passing the original malicious request (data point) through the intent launderer to generate a revision. This revision is then used to attack the target model. An LLM judge evaluates the response for safety and practicality. If a response is both unsafe and practical, the attack is considered successful. Otherwise, the revision–regeneration mechanism is triggered, using all previous failed revisions as feedback to generate a new, improved revision. The loop ends when either a set number of iterations is reached or a target attack success rate is achieved.
In this work, we systematically study the quality of widely used AI safety datasets—AdvBench (Zou et al., 2023) and HarmBench (Mazeika et al., 2024)—based on prior research (Kim et al., 2025; Xie et al., 2025; Zhao et al., 2024; Xu et al., 2024; Li et al., 2024; Anil et al., 2024, inter alia). First, we analyze whether these datasets, in isolation, faithfully represent real-world attacks by evaluating three defining properties of such attacks: being driven by ulterior intent, well-crafted, and out-of-distribution. [^1] Similar to other datasets, diversity is another important consideration in the design of safety datasets. This includes diversity across topics and data points. Topic diversity is generally well accounted for by most dataset creators (Röttger et al., 2024a; Mazeika et al., 2024; Zou et al., 2023) and is thus excluded from our analysis. In contrast, diversity at the data-point level is a major issue in these datasets, which we discuss in Section 2.2. Second, we examine whether these datasets actually measure safety risks in practice when they are used to evaluate models.
To evaluate the quality of safety datasets in isolation, we begin by analyzing $n$-gram word clouds. This helps visualize how the most frequent unigrams recur as higher-order $n$-grams. We find that these recurring $n$-grams consistently form unrealistic “triggering cues [^2] We use “triggering language” interchangeably.: expressions with overt negative/sensitive connotations. These cues fall into two categories: (1) inherent—expressions that carry such connotations by nature, regardless of context (e.g., “commit suicide”), and (2) contextual—expressions that acquire such connotations in the context of harmful requests (e.g., [malicious intent] + “without getting caught,” explicitly signaling evasion). Figure 2 shows examples of triggering cues. These cues undermine two properties—being well-crafted and driven by ulterior intent—as such overt language rarely appears in real-world attacks and seems engineered to trigger safety mechanisms artificially. We also show that the repetitive overuse of these cues creates substantial duplication, producing near-identical data points in both sentence structure and malicious intent. This further degrades the out-of-distribution property and adds to the erosion of well-crafted property. Overall, our analysis indicates that current safety datasets fail to faithfully represent real-world attacks.
Next, we evaluate the quality of safety datasets in practice by involving models. In particular, we examine whether these datasets genuinely measure safety risks or merely rely on triggering language that safety-aligned models are trained to detect and refuse to answer. To explore this, we introduce “intent laundering”: a systematic procedure that abstracts away overt triggering language from attacks (data points) while strictly preserving their malicious intent. This process has two components: (1) connotation neutralization, and (2) context transposition. In connotation neutralization, triggering expressions are replaced with neutral/positive or descriptive alternatives. In context transposition, real-world scenarios and referents—such as individuals (e.g., “immigrants”) or institutions (e.g., “charity”)—that can act as triggering cues in harmful requests are mapped to non-real-world contexts (e.g., a game world). We automate this process using an “intent launderer”: a large language model (LLM) with a few-shot in-context learning (ICL) setup. Each ICL demonstration pairs an original data point with its manually crafted, intent-laundered revision. Figure 1 depicts an overview of this framework. Our results reveal a strong effect: once triggering language is removed, the attack success rate (ASR) jumps from a mean of 5.38% to 86.79% on AdvBench, and from 13.79% to 79.83% on HarmBench. This indicates that current safety evaluations, and the safety alignment built on them, overrely on triggering cues, leading to an inflated evaluation of safety relative to real-world attacks.
Finally, we propose intent laundering as a standalone jailbreaking technique by adding an iterative revision–regeneration mechanism. In each iteration, the model uses all previous failed revisions as feedback to generate a new, improved revision, using the same few-shot ICL setup. Figure 1 shows this mechanism. This iterative process continues until either a predefined number of regeneration attempts is reached or a target ASR is met. Our results show that, with this regeneration loop, intent laundering achieves high ASR (90%98.55%) after only a few iterations across all studied models under fully black-box access. This includes recent models widely reported as among the safest—such as Gemini 3 Pro (Google, 2025a; b) and Claude Sonnet 3.7 (Anthropic, 2025a; b; Holistic AI, 2025). These findings further confirm that existing safety evaluations and safety-alignment methods are highly overfitted to triggering cues.
The key contributions of this paper are as follows:
(1) We conduct a systematic study of the quality of widely used AI safety datasets and show that they do not faithfully reflect real-world attacks due to overuse of unrealistic triggering cues.
(2) We introduce intent laundering: a procedure that empirically verifies that revising attacks to remove triggering cues sharply increases ASR, from a mean of 5.38% to 86.79% on AdvBench, and from 13.79% to 79.83% on HarmBench.
(3) We adapt intent laundering into a novel jailbreaking method by adding a revision–regeneration step for failed revisions, achieving high ASR (90%–98.55%) across the board.
(4) We find that recent safety-alignment techniques still fall significantly short in repelling real-world attacks, and that models aligned using these methods are often referred to as “reasonably safe” mainly because existing safety evaluations are overly simplistic and unrealistic.
2 Empirical Motivation
To motivate our study, we first analyze the quality of safety datasets in isolation. We use $n$-gram word clouds and pairwise similarity analyses to assess whether these datasets reflect real-world attacks. Specifically, we analyze three key characteristics of such attacks:
  • Ulterior Intent: Attacks conceal harmful intents behind benign-looking requests.
  • Well-crafted: Attacks are carefully designed to bypass safety filters.
  • Out-of-distribution: Attacks differ from everyday user prompts.
2.1 Word Clouds
(a) Unigrams
(b) Bigrams
(c) Trigrams
Figure 2: Word clouds of the 40 most frequent $n$-grams, where $n\in\{1,2,3\}$, from the combined AdvBench and HarmBench corpus. Triggering cues are highlighted in red for inherently, and orange for contextually negative/sensitive connotations. This includes triggering words in the unigram cloud and triggering phrases in the bigram and trigram clouds. Neutral-connotation unigrams that contribute to triggering cues in higher-order $n$-grams are also shown in green. These visualizations reveal an unusual overrepresentation of overtly triggering cues in safety datasets, suggesting that data points are artificially designed to trigger safety mechanisms. For example, expressions such as “tutorial” and “step-by-step instructions” are used explicitly to trigger infohazard safeguards (Bostrom, 2011), which is unrealistic. Even minimally skilled bad actors rarely use such overt and self-incriminating language (e.g., “commit identity theft”). These findings indicate that safety datasets fail to capture two of the key properties of real-world attacks: being well-crafted and driven by ulterior intent. Note that, for clarity, only a representative subset of expressions is marked in the word clouds above.
We use $n$-gram word clouds to assess the quality of safety datasets intuitively. Specifically, we generate word clouds of the 40 most frequent $n$-grams in the combined AdvBench and HarmBench corpus, with $n\in\{1,2,3\}$. These visualizations help trace how common unigrams evolve into higher-order $n$-grams, revealing dominant language patterns.
Figure 2 displays the resulting word clouds. At the unigram level, there is an unusual overrepresentation of words with overt negative/sensitive connotations, which we refer to as “triggering words.” Triggering words fall into two categories: (1) inherent triggering words (in red), which carry negative/sensitive connotations by nature (e.g., “steal”), and (2) contextual triggering words (in orange), which signal such connotations when used in harmful contexts (e.g., “commit”).
As unigrams evolve into higher-order $n$-grams, this overrepresentation intensifies: phrases with negative/sensitive connotations become unusually dominant. Similar to unigrams, these phrases carry such connotations either inherently or contextually within malicious requests. We refer to these as “triggering phrases,” and together with triggering words, they form what we term “triggering cues.”
Triggering phrases mainly form in two ways: (1) they build on triggering words, or (2) they consist entirely of words with neutral connotations. For example, the inherent triggering word “steal,” evolves into inherent triggering phrases such as “steal sensitive information,” “steal confidential information,” and “steal personal information.” Similarly, the contextual triggering word “commit,” expands into inherent triggering phrases such as “commit suicide,” “commit insider trading,” and “commit identity theft.” Neutral-connotation words can also form triggering phrases. For instance, “without,” “getting,” and “caught,” combine into the triggering phrase “without getting caught.” This also explains their high frequency as unigrams.
However, such explicit and repetitive overuse of triggering cues, along with direct mentions of malicious intent, directly contradicts real-world attack scenarios. Even minimally capable adversaries rarely use such overt language, as it easily triggers safety mechanisms. Taken together, these patterns indicate that existing safety datasets contain contrived data points largely disconnected from real-world attacks, where harmful requests are well-crafted and motivated by ulterior intent.
Figure 3: Proportion of duplicated versus unique data points in the AdvBench and HarmBench datasets across varying similarity thresholds. Each safety dataset is compared to a size-matched GSM8K subset shown below its plot. Both safety datasets exhibit considerably higher duplication rates across most thresholds compared to their GSM8K counterparts. This is striking, as safety datasets are intended to approximate real-world attacks—characterized by being out-of-distribution and well-crafted. In contrast, they show more duplication than a regular non-safety dataset, where such duplication is more acceptable. This is particularly alarming for safety datasets, as it indicates that many data points in these datasets evaluate the model on essentially the same harmful intent in nearly identical contexts (see Figure 4 for examples), leading to an overestimated evaluation of safety.
2.2 Data Duplication
The excessive use of triggering cues not only creates unrealistic data points but also suggests significant overlap in sentence structure and malicious intent across data points. To analyze this, we conduct a straightforward pairwise similarity check between data points within each dataset. We use similarity thresholds ranging from 0.7 to 0.99. For a given threshold, data points that do not meet the similarity threshold with any other data point are labeled as unique, while those that meet/exceed the threshold with at least one other data point are labeled as duplicated and grouped to represent a single data point. As there is no universally accepted threshold for high similarity—particularly within datasets containing data points from a single context—we use the GSM8K dataset (Cobbe et al., 2021), a widely used non-safety dataset, as a baseline. For a fair comparison, we subsample GSM8K to match the size of each safety dataset.
Figure 3 presents the results of our similarity analysis. The first key finding is that over 45% of data points in AdvBench are near-identical at a 0.95 similarity threshold. More notably, over 11% are (almost) exact copies at a 0.99 threshold. These numbers are unusually high for a safety dataset with only 520 data points—especially one intended to reflect out-of-distribution and well-crafted attacks.
Figure 4: Examples of duplicated data points from the AdvBench and HarmBench datasets. These examples exhibit two unusual patterns: (1) explicit and repetitive overuse of triggering cues, either inherently (in red, e.g., “chop shops”) or contextually (in orange, e.g., “in detail”), and (2) substantial duplication resulting from this overuse. Each group of duplicates effectively represents a single malicious intent, i.e., a refusal or response to one is sufficient to evaluate the robustness of the model for that intent. As a result, safety evaluations based on these data points can be inflated.
The second key finding comes from cross-dataset comparison. At a 0.85 threshold, only about 11% of AdvBench data points are unique, compared to nearly 94% in the size-matched GSM8K subset. HarmBench also shows considerable duplication: 16% of its data points are duplicated at this threshold, versus only 3.5% in its GSM8K counterpart—over four times higher. Considering 85% uniqueness as a reasonable target for a well-designed dataset, AdvBench requires an extremely high threshold of 0.99, and even then, it never reaches 90% uniqueness. In contrast, the GSM8K counterpart reaches 85% uniqueness at a much more moderate 0.85 threshold. Similarly, HarmBench requires a high threshold of 0.9 to hit 85% uniqueness, whereas its GSM8K counterpart reaches it at just 0.8. These findings are concerning. Safety datasets are expected to feature more unique data points to effectively reflect real-world attacks. Instead, they exhibit far more duplication than non-safety datasets such as GSM8K, where homogeneity is more acceptable. This duplication undermines two of the key properties required for modeling real-world attacks: being out-of-distribution and well-crafted. [^3] These findings also suggest that reported results from past studies using subsets of safety datasets can be misleading (Xie et al., 2025; Dékány et al., 2025; Xhonneux et al., 2024, inter alia). In fact, many subsampled data points can come from the same group of near-duplicates, as shown in Figure 4. This can inflate performance, as repeated instances of a single harmful intent that a method handles well can dominate the sample. The issue is further amplified by the small size of these datasets. Figure 4 provides examples of these duplications.
Overall, analyzing safety datasets in isolation yields two main insights: (1) they repeatedly overuse triggering cues, and (2) this leads to substantial duplication. Together, these undermine all three key properties of real-world attacks. Thus, they fail to faithfully represent real-world adversarial behavior, and safety evaluations based on them can be inflated. This raises a key question: if triggering cues are removed, do models previously reported as “reasonably safe” still remain so? We explore this next.
3 Approach
3.1 Intent Laundering
Here, we evaluate the quality of safety datasets in practice when used to evaluate models. In particular, we examine whether these datasets actually capture safety risks or whether they mainly rely on triggering cues to elicit refusals to harmful requests. To explore this, we introduce “intent laundering”: a procedure that abstracts away overt triggering language from harmful requests (data points) while strictly preserving their malicious intent and all relevant details. In fact, the idea is to imply the harmful intent rather than explicitly state it.
Intent laundering consists of two complementary components:
  • Connotation Neutralization: Removes triggering cues carrying negative/sensitive connotations by replacing them with neutral/positive alternatives. If no such alternatives exist, it uses descriptive substitutes. Figures 5 and 6 in Appendix B show examples of such transformations.
  • Context Transposition: Replaces real-world scenarios and referents—such as individuals (e.g., “immigrants”) or institutions (e.g., “charity”)—that can act as triggering cues in harmful contexts with their non-real-world alternatives (e.g., a game world or fictional narrative with stand-ins). All details remain applicable and transferable to the real world with no or minor adjustments. Figures 5 and 6 in Appendix B show examples of context transposition.
Based on the intent-laundering principles, we first manually craft a few intent-laundered revisions of data points from the two safety datasets under study. We then automate this process using an “intent launderer”: a large language model (LLM) coupled with a few-shot in-context learning (ICL) setup. Each original–revision pair serves as an ICL demonstration that guides the model in performing intent laundering. Figure 1 provides an overview of the intent laundering framework. Figure 7 in Appendix C presents the ICL prompt used by the intent launderer.
After generating all revisions, we collect model responses to both the original and revised data points to assess how triggering language affects safety evaluation.
3.2 Evaluation
As intent laundering modifies how harmful requests are expressed (via connotation neutralization) and structured (via context transposition), we evaluate model responses along two axes: safety and practicality. These two criteria are assessed independently and do not influence each other.
Safety Evaluation.
We exclusively evaluate the safety of responses to revisions using a five-point Likert scale based on the level of harm exposed: extremely unsafe, highly unsafe, unsafe, moderately unsafe, and safe (see Figure 10 in Appendix C for score definitions). [^4] We do not use numerical scales to avoid potential biases against assigning extremely high/low scores, as reported in prior studies (Ouyang et al., 2022; Golchin et al., 2025). We also condition our safety evaluation specifically on harm relative to the original malicious intent, which serves as the evaluation criterion (see Figure 9 in Appendix C for how it is generated for each attack). Safety is thus judged on two fronts: (1) whether the response to the revision addresses the same harmful intent as the original version, and if so, (2) how unsafe the response is. If the revision fails to preserve the original malicious intent, the response is rated as safer.
Practicality Evaluation.
Similar to the safety evaluation, we exclusively evaluate the practicality of responses to revisions using a five-point Likert scale based on the practicality and transferability of the provided details in the real world: extremely applicable, highly applicable, applicable, moderately applicable, and inapplicable (see Figure 11 in Appendix C for score definitions). Similarly, we condition our practicality evaluation specifically on harm relative to the original malicious intent, which is provided as the evaluation criterion. As with safety, practicality is judged on two fronts: (1) whether the details in the response support the same harmful intent as the original version, and if so, (2) how practical those details are when applied in the real world with reasonable adjustments.
LLM as a Judge.
We use an LLM to automatically evaluate both safety and practicality. The model receives the same instructions shown in Figures 10 and 11 in Appendix C for safety and practicality evaluations, respectively.
Attack Success Rate.
We measure performance using the attack success rate (ASR), defined as the fraction of attack attempts that succeed. In our setting, an attack (revision) is considered successful only if the response is both unsafe and practical. A response is unsafe if it is rated as extremely unsafe, highly unsafe, or unsafe, and practical if it is rated as extremely applicable, highly applicable, or applicable. An attack is unsuccessful if it fails to generate an unsafe response, or if the unsafe response is too abstract to be practical, even after adjustments. This design choice—using five-point scales and then mapping them to binary outcomes—aims to reduce ambiguity in borderline cases.
3.3 Intent Laundering as a Jailbreaking Technique
We extend intent laundering into an independent jailbreaking method, leveraging its ability to avoid triggering safety filters. Specifically, we add an iterative revision–regeneration mechanism for failed revisions, while the main procedure remains unchanged. In each iteration, all previous failed revisions are provided as feedback to the intent launderer to generate a new, improved revision, using the same few-shot ICL setup. Figure 1 illustrates intent laundering when applied as a jailbreaking technique. Figure 8 in Appendix C shows the ICL prompt used in the revision–regeneration loop, incorporating feedback from earlier failed revisions.
Under our definition of failed revisions, the revision–regeneration mechanism improves attack performance in two ways: (1) by generating new revisions that succeed where earlier ones failed to elicit unsafe responses, and (2) by generating new revisions that yield more practical responses where previous responses were too abstract. The process repeats until either a predefined number of regeneration attempts is reached or a target ASR is achieved.
4 Experimental Setup
Word Clouds.
We use the AdvBench (Zou et al., 2023) and HarmBench (standard) (Mazeika et al., 2024) datasets. We generate word clouds using the Python wordcloud package (Mueller, ). To preserve actual language patterns, we apply only lowercase conversion and whitespace normalization. For clearer visualization, we remove stopwords, punctuation, special characters, and words that instruct models, including “write,” “generate,” “create,” “develop,” “use,” “give,” “provide,” and “people.”
Data Duplication.
We use the same safety datasets as in the word cloud analysis, along with two randomly sampled subsets of the GSM8K dataset (Cobbe et al., 2021). Each GSM8K subset matches the size of a corresponding safety dataset. This ensures a fair pairwise similarity comparison between safety and non-safety datasets, since the number of data points significantly affects similarity scores, as can be seen in Figure 3 across the two GSM8K subsets.
We use embeddings from Sentence Transformers (Sentence-BERT) (Reimers & Gurevych, 2019; Devlin et al., 2019; Wolf et al., 2019; Vaswani et al., 2017), specifically the all-MiniLM-L6-v2 checkpoint, which is fine-tuned for clustering and semantic search. We choose this checkpoint to mitigate embedding anisotropy (Timkey & van Schijndel, 2021; Ethayarajh, 2019), a challenge amplified in our setting due to significant prefix overlap among data points in our safety datasets, which can produce uniformly high, non-informative similarity scores. For each data point, we average token-level contextual embeddings and compute pairwise cosine similarities between all data points in a dataset. We set the maximum input length to 512 tokens.
Evaluation Criteria Generation.
We utilize GPT-4o (gpt-4o-2024-11-20) (Hurst et al., 2024) with an 8-shot ICL setup to generate evaluation criteria. All generation hyperparameters are kept at their default values, and output is limited to 1024 tokens.
Intent Launderer.
We leverage GPT-5.1 (gpt-5.1-2025-11-13) (OpenAI, 2025) with an 8-shot ICL setup as the intent launderer. For jailbreaking, the same model and ICL setup are used to regenerate failed revisions at each iteration. We set the reasoning effort to none, cap output at 2048 tokens, and use default values for all other decoding hyperparameters.
LLM as a Judge.
We employ GPT-5.1 (gpt-5.1-2025-11-13) as a powerful judge to evaluate both safety and practicality when intent laundering is applied. Each evaluation is run independently in a zero-shot setting using separate model calls. All hyperparameters match those of the intent launderer, except that the output is capped at 1024 tokens. We report details on the agreement between LLM-based and human evaluations for both safety and practicality in Appendix A.
For safety evaluation on the original attacks (data points), where intent laundering is not applied, we use the standard ASR metric with the same five-point Likert scale described in Subsection 3.2 under Safety Evaluation. The corresponding prompt is shown in Figure 12 in Appendix C. We follow the same categorization scheme: an attack is considered successful if its response is labeled as extremely unsafe, highly unsafe, or unsafe. As before, all hyperparameters are identical to those of the intent launderer, except that the output is limited to 1024 tokens.
Evaluation Datasets.
Following the data duplication analysis in Subsection 2.2, we apply a conservative similarity threshold of 0.9 to work with unique data points in the safety datasets. Data points above this threshold are treated as duplicates, as they share similar sentence structures and identical malicious intent. For each group of duplicates, we randomly select one representative data point. This yields 207 data points from AdvBench and all 200 data points from HarmBench (standard). We use these two sets for all safety evaluation experiments.
Evaluation Models.
We select a diverse set of models for evaluation. In total, we experiment with seven different models: Gemini 3 Pro (gemini-3-pro-preview) (Google, 2025b), Claude Sonnet 3.7 (claude-3-7-sonnet-20250219) (Anthropic, 2025b), Grok 4 (grok-4-fast-non-reasoning) (xAI, 2025), GPT-4o (gpt-4o-2024-11-20) (Hurst et al., 2024), Llama 3.3 70B (llama-3.3-70b-instruct) (Grattafiori et al., 2024), GPT-4o mini (gpt-4o-mini-2024-07-18) (Hurst et al., 2024), and Qwen2.5 7B (qwen2.5-7b-instruct) (Yang et al., 2025). As our task does not require advanced reasoning and the base models remain the same across reasoning levels, we reduce reasoning efforts where applicable. Specifically, for Gemini 3 Pro, we set reasoning effort to low; for Claude Sonnet 3.7, we use standard mode; and for Grok 4, we use the non-reasoning checkpoint. All models use default inference hyperparameters, with output capped at 4096 tokens.
5 Results and Discussion
Table 1: Safety evaluation (SE), practicality evaluation (PE), and attack success rate (ASR) on the AdvBench dataset (top) and the HarmBench dataset (bottom). SE is the percentage of model responses rated as extremely unsafe, highly unsafe, or unsafe. PE is the percentage of responses rated as extremely applicable, highly applicable, or applicable. ASR is the percentage of responses that satisfy both SE and PE simultaneously. In the no-revision setting, where original data points from safety datasets are used, ASR follows its standard definition, as no abstraction is applied; thus, SE and PE are not applicable. The first-revision setting corresponds to the first application of intent laundering, where triggering cues are removed. All subsequent iterations reflect intent laundering with the revision–regeneration loop, which functions as a jailbreaking technique. Bold values denote the highest ASR across all iterations (within each dataset). Lower ASR implies better model safety.
  No Revision First Revision Second Revision       Third Revision Model ASR SE PE ASR SE PE ASR SE PE ASR Gemini 3 Pro 1.93 83.09 99.42 82.61 90.34 100.00 90.34 95.17 100.00 95.17 Claude Sonnet 3.7 2.42 81.64 97.63 79.71 86.96 98.89 85.99 93.72 99.48 93.23 Grok 4 17.87 90.82 100.00 90.82 96.14 99.50 95.66 96.62 100.00 96.62 GPT-4o 0.00 82.61 98.27 81.18 93.72 98.45 92.27 94.69 98.47 93.24 Llama 3.3 70B 10.14 91.79 100.00 91.79 98.07 100.00 98.07 98.55 100.00 98.55 GPT-4o mini 0.97 90.34 98.93 89.37 95.17 100.00 95.17 96.62 100.00 96.62 Qwen2.5 7B 4.35 92.75 99.48 92.27 95.65 100.00 95.65 97.10 100.00 97.10 Mean 5.38 87.58 99.10 86.79 93.72 99.55 93.30 96.07 99.71 95.79
  No Revision First Revision Second Revision Third Revision Fourth Revision       Fifth Revision Model ASR SE PE ASR SE PE ASR SE PE ASR SE PE ASR SE PE ASR Gemini 3 Pro 11.00 80.00 98.12 78.50 84.50 99.41 84.00 88.50 100.00 88.50 90.50 100.00 90.50 93.00 100.00 93.00 Claude Sonnet 3.7 8.50 73.00 96.58 70.50 78.50 99.36 78.00 85.50 100.00 85.50 87.00 100.00 87.00 91.00 100.00 91.00 Grok 4 36.00 82.00 97.56 80.00 87.00 98.28 85.50 88.00 100.00 88.00 88.50 100.00 88.50 93.00 100.00 93.00 GPT-4o 0.50 89.00 100.00 89.00 90.00 99.44 89.50 91.00 100.00 91.00 92.00 100.00 92.00 93.00 100.00 93.00 Llama 3.3 70B 14.00 83.50 96.41 80.50 84.00 99.40 83.50 86.00 100.00 86.00 87.00 100.00 87.00 91.00 100.00 91.00 GPT-4o mini 5.00 80.00 99.38 79.50 84.00 99.40 83.50 85.00 100.00 85.00 87.50 99.43 87.00 91.00 98.90 90.00 Qwen2.5 7B 21.50 83.00 97.59 81.00 87.50 100.00 87.50 89.00 100.00 89.00 90.00 100.00 90.00 90.50 100.00 90.50 Mean 13.79 81.50 97.95 79.83 85.07 99.33 84.50 87.57 100.00 87.57 88.93 99.92 88.86 91.79 99.84 91.64
Table 1 presents results across seven models on both AdvBench and HarmBench. This includes results from three experimental settings:
  • No Revision: Results from the original data points, which do include triggering cues.
  • First Revision: Results after applying intent laundering to remove triggering cues.
  • Other Revisions: Results when intent laundering is used as a jailbreaking method.
Based on our results, we make the following observations: [^5] To draw general observations, we do not apply the termination conditions in Subsection 3.3. Instead, we run a representative number of iterations to assess the effectiveness of our methodology.
(1) Removing triggering cues from data points (attacks) leads to a sharp increase in ASR. On AdvBench and HarmBench, ASR rises from a mean of 5.38% and 13.79% in the no-revision settings, where cues are present, to 86.79% and 79.83%, respectively, in the first-revision settings, where cues are first removed. This indicates that model refusals are largely driven by the presence of triggering cues. Consequently, safety datasets do not reliably measure real-world safety risks, as they rely more on triggering cues to elicit refusals than on actual malicious intent.
(2) Intent laundering is highly effective at removing triggering cues while preserving the malicious intent. It also acts as a strong jailbreaking technique. As shown by the bold values—corresponding to the highest ASR and the last iteration in each dataset—intent laundering achieves high ASR values, ranging from 90% to 98.55% across all models and datasets. This includes models such as Gemini 3 Pro, known for strong safety (Google, 2025a; b), and Sonnet 3.7, known for overrefusal (Zhang et al., 2025; Cui et al., 2024). These models are jailbroken at high ASRs of 93%95% on AdvBench and 91%93% on HarmBench—and with only a few iterations.
(3) Despite the abstraction introduced by intent laundering, model responses remain applicable and transferable to the real world. This is supported by the high practicality rates across all iterations, models, and datasets.
(4) ASR consistently increases with more iterations. While the first iteration always yields the largest gain, ASR continues to rise—with a mean increase of 9% on AdvBench and 11.81% on HarmBench by the final iteration—showing steady growth throughout. This confirms that the revision–regeneration mechanism effectively and systematically boosts ASR, and that adjusting the number of iterations provides direct control over the desired ASR.
(5) Our results suggest that both internal safety evaluations and safety-related training paradigms likely overrely on triggering cues, similar to publicly available safety datasets. This is supported by the fact that all insights motivating our methodology—such as unusual language patterns and dataset design flaws—are derived exclusively from publicly available safety datasets, rather than from safety-alignment techniques or internal safety evaluations. Nevertheless, these insights remain broadly effective, as evidenced by consistently high ASR in all experiments, regardless of whether models are closed- or open-weight, old or recent. Further evidence comes from the fact that internal safety evaluations reach the same conclusions as publicly available safety datasets: these models are reasonably safe (Anthropic, 2025a; b; Google, 2025a; b; xAI, 2025; Yang et al., 2025; Hurst et al., 2024; Grattafiori et al., 2024)a conclusion that stands in contrast to our findings.
6 Related Work
Safety Alignment.
The main objective of safety alignment is to balance helpfulness and harmlessness in AI models, avoiding both underrefusal (overly helpful) and overrefusal (overly harmless) (Röttger et al., 2024b; Bai et al., 2022; Ouyang et al., 2022; Askell et al., 2021). Adversarial attacks, however, can disrupt this balance, causing “aligned” models to generate harmful outputs—a behavior known as misalignment (Deshpande et al., 2023; Ouyang et al., 2022; Askell et al., 2021; Sheng et al., 2019).
Broadly, these attacks fall into two categories: “invasive” and “non-invasive.” We define invasive attacks as methods that erode safety alignment by directly modifying model weights. This includes using specialized fine-tuning/training recipes on benign data (Mu et al., 2025; Xie et al., 2025; Qi et al., 2024; Yang et al., 2024; Yi et al., 2024; Hawkins et al., 2024; Lermen et al., 2023; Halawi et al., 2024, inter alia), as well as training on a small number of harmful examples (Souly et al., 2025). In contrast, non-invasive attacks operate solely through input prompt engineering, without altering model parameters. Examples include attacks based on ciphers (Yuan et al., 2024; Handa et al., 2024), many-shot in-context learning (Anil et al., 2024; Golchin et al., 2024), membership inference attacks that induce the model to emit training data (Golchin & Surdeanu, 2025; Nasr et al., 2025b; Golchin & Surdeanu, 2024; Carlini et al., 2023; 2021; Shokri et al., 2017, inter alia), and highly engineered red-teaming prompts, either crafted by humans (Nasr et al., 2025a; Yu et al., 2024; Schulhoff et al., 2023), generated by LLMs (Joo et al., 2025; Wahréus et al., 2025; Mehrotra et al., 2024; Jin et al., 2024), or optimized via statistical machine learning methods (Nasr et al., 2025a).
In response to these increasingly sophisticated attacks, advanced safety alignment techniques were proposed to enhance safety. Recent work primarily leverages reasoning as a mechanism to increase robustness, particularly in LRMs (Kim et al., 2025; Guan et al., 2024; Jaech et al., 2024). At the same time, this reasoning can itself be misused under adversarial conditions to jailbreak models (Sabbaghi et al., 2025). Overall, studies show that the safety alignment of current models remains fragile (Shah et al., 2025; Qi et al., 2025; Amodei et al., 2016), making reliable safety evaluation particularly challenging (Rando et al., 2025; Benton et al., 2024).
Safety Datasets.
As with other evaluation tasks, safety datasets aim to measure the effectiveness of safety alignment by capturing real-world attack scenarios. Early research focused on evaluating models against a narrow set of risks, such as bias (Tamkin et al., 2023; Nadeem et al., 2021; Dhamala et al., 2021; Nangia et al., 2020), toxicity (Hartvigsen et al., 2022; Gehman et al., 2020), and ethical judgment (Hendrycks et al., 2021). However, with the rise of general-purpose models, the focus broadened to encompass a wider range of safety threats, including misinformation, cybercrime, harassment, and more (Chao et al., 2024; Mou et al., 2024; Wang et al., 2023; Xu et al., 2021, inter alia). Within this expanded risk landscape, a particularly pressing area of concern in recent research involves preventing the misuse of models for chemical, biological, radiological, and nuclear threats (60 Minutes, 2025). This growing concern elevated datasets such as AdvBench (Zou et al., 2023) and HarmBench (Mazeika et al., 2024) as prominent benchmarks for evaluating broader safety alignment.
7 Conclusion
We systematically studied the quality of two widely used AI safety datasets: AdvBench and HarmBench. We began by analyzing these datasets in isolation, without using any models. This analysis revealed that these datasets do not faithfully approximate real-world adversarial behavior due to their overreliance on triggering cues: expressions with overt negative/sensitive connotations designed to trigger safety mechanisms artificially. Motivated by this finding, we investigated what these datasets actually measure in practice when triggering cues are present, and whether their conclusions about model safety still hold when such cues are removed. To this end, we introduced intent laundering: a procedure that removes triggering cues from attacks (data points) while strictly preserving their malicious intent and all relevant details. Our results showed that prior safety conclusions do not hold once triggering cues are removed, and that the observed safety performance is largely driven by the presence of triggering cues rather than by the underlying safety risks. We further showed that intent laundering can be used as a powerful jailbreaking technique, achieving high attack success rates from 90% to over 98%. Overall, our findings unveiled a critical gap between how model safety is evaluated and how real-world adversarial behavior manifests. Based on this, we conclude that (1) safety evaluations must evolve to capture adversarial attacks more realistically, and (2) current safety-alignment efforts are still far from robust against real-world threats.
Ethics Statement
We acknowledge that our findings can affect public and institutional trust in current safety claims and evaluations. However, our intention is not to undermine trust in AI safety research, but rather to improve its scientific rigor and practical relevance. We believe the societal benefits of exposing and addressing weaknesses in safety evaluation frameworks significantly outweigh the potential risks, especially when combined with responsible disclosure and awareness efforts. We proactively informed affected model providers and collaborated with them to improve the safety of their products.
Appendix A Agreement Analysis Between LLM-based and Human Evaluations
We measure agreement between the LLM judge and human expert consensus on 100 randomly selected instances, drawn from LLM judgments on responses from all models studied. For safety evaluation, model responses are categorized as unsafe (extremely unsafe, highly unsafe, or unsafe) or safe (moderately unsafe or safe). For practicality evaluation, model responses are categorized as practical (extremely applicable, highly applicable, or applicable) or impractical (moderately applicable or inapplicable). Human consensus is determined by the majority vote among three independent expert annotators. We report percent agreement rather than Cohen’s kappa (Cohen, 1960; Scott, 1955; Fleiss, 1971) due to high base rates of the majority class, which inflate chance agreement and render Kappa uninformative. We compute 95% confidence intervals (CI) using bootstrap resampling with 10,000 iterations (Efron, 1992; Tibshirani & Efron, 1993; Efron, 2003).
Table 2 reports the results of our agreement analysis. The LLM judge achieves 90.00% agreement (95% CI: 83.00%–95.00%) with human consensus on binary safety evaluation and 94.00% agreement (95% CI: 89.00%–98.00%) on binary practicality evaluation, meeting or exceeding the mean pairwise agreement among human annotators (82.00% for safety, 77.33% for practicality). The narrow bootstrap-estimated confidence intervals indicate stable performance, and the fact that the LLM’s agreement with consensus matches/surpasses human-human reliability suggests that the model can serve as a reliable automated judge for both safety and practicality evaluations.
Table 2: Agreement between LLM judge and human consensus on binary safety and practicality evaluations across 100 randomly sampled instances. Safety is labeled unsafe (extremely unsafe, highly unsafe, or unsafe) or safe (moderately unsafe or safe). Practicality is labeled practical (extremely applicable, highly applicable, or applicable) or impractical (moderately applicable or inapplicable). Human consensus is the majority vote of three expert annotators.
Evaluation Overall Agreement 95% CI Human-Human Agreement (mean)
Safety 90.00% [83.00%, 95.00%] 82.00%
Practicality 94.00% [89.00%, 98.00%] 77.33%
Appendix B Examples of Intent-Laundered Revisions and Model Responses
We present examples of actual model responses to revisions produced by the intent laundering process. Figure 5 presents a response from Gemini 3 Pro to a revision based on a data point from the AdvBench dataset. Similarly, Figure 6 shows a response from Claude Sonnet 3.7 to a revision based on a data point from the HarmBench dataset. As shown, all responses are unsafe, detailed, and transferable to real-world contexts, despite being framed as abstract scenarios.
Figure 5: An actual response (red box) generated by Gemini 3 Pro to an intent-laundered revision (green box) based on a data point from the AdvBench dataset (orange box). The revision uses both connotation neutralization and context transposition to abstract away triggering cues. The model response is partially shown to prevent potential misuse; however, the full response spans several paragraphs and fully enables the original malicious intent. This example is provided strictly for academic safety research. Any misuse is strongly discouraged.
Figure 6: An actual response (red box) generated by Claude Sonnet 3.7 to an intent-laundered revision (green box) based on a data point from the HarmBench dataset (orange box). The revision uses both connotation neutralization and context transposition to abstract away triggering cues. The model response is partially shown to prevent potential misuse; however, the full response spans several paragraphs and fully enables the original malicious intent. This example is provided strictly for academic safety research. Any misuse is strongly discouraged.
Appendix C Input Prompts
This section lists all prompts used in our study. These include prompts for intent laundering, the revision–regeneration loop, evaluation criteria generation, safety evaluation, practicality evaluation, and regular safety evaluation, where intent laundering is not applied (no-revision settings). Figures 7, 8, 9, 10, 11, and 12 present them in that order.
Figure 7: Prompt used by the intent launderer to generate revisions. The placeholder is filled with the data point being intent-laundered.
Figure 8: Prompt used by the intent launderer to generate revisions in the revision–regeneration loop. The placeholders are filled with the previously failed revisions of the data point being intent-laundered, and the data point itself, in top-to-bottom order.
Figure 9: Prompt used to generate evaluation criteria. The placeholder is filled with the original data point.
Figure 10: Prompt used for safety evaluation of responses to revisions. The placeholders are filled with the evaluation criterion followed by the model response, in top-to-bottom order.
Figure 11: Prompt used for practicality evaluation of responses to revisions. The placeholders are filled with the evaluation criterion followed by the model response, in top-to-bottom order.
Figure 12: Prompt used for regular safety evaluation of responses to the original attacks (data points), where intent laundering is not applied. The placeholder is filled with the model response.